Andrea, a self-employed individual, wishes to accumulate a retirement fund of $650,000. How much should she deposit each month into her retirement account, which pays interest at a rate of 2.5%/year compounded monthly, to reach her goal upon retirement 25 years from now? (Round your answer to the nearest cent.) $
Answer #1

[(1+r) -1 FV of Annuity = P P= Periodic Payment r=rate per period n = number of periods

n = 25 * 12 = 300 months

FV = $650,000

r = 2.5%/12 = 0.2083% (monthly)

650,000 = P* (1 + 0.002083) 300 - 1 0,002083

650,000 = P * 416.1752

P = $1,561.8